# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 6.1 Reliability of the a posteriori error estimator
$\begin{array}[]{ll}\displaystyle|\mathcal{R}_{1}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))|\leq C_{4}\gamma\|p-p_{h}||_{0,\Omega}\|{\boldsymbol{\chi}}\|_{1,\Omega}\\[8.61108pt] \displaystyle+C_{4}\biggr{(}\,\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}||\underline{\mathbf{curl}}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})||_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Omega)}h_{E}||[\![(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}]\!]||_{0,E}^{2}\\[8.61108pt] \displaystyle\quad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}||(C^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}-\dfrac{\mathrm{d}\boldsymbol{u}_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}||_{0,E}^{2}\biggr{)}^{1/2}\|{\boldsymbol{\chi}}\|_{1,\Omega}.\end{array}$
Proof 6.8.
Similarly to Lemma 6.5, adding ±$\pm\,c\,(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}),p_{h})$, we have
@@
(6.5) $\begin{array}[]{ll}\mathcal{R}_{1}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))\\ \quad\displaystyle=H(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))-c(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}),p)-a(\boldsymbol{\sigma}_{h},\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))-b(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}),(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))\\ \displaystyle\quad=\langle(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))\\ \displaystyle\quad-\int_{\Omega}\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}:\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})-\int_{\Omega}\boldsymbol{\rho}_{h}:\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p_{h}\,\mathop{\rm tr}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})).\end{array}$
@@
Then, applying a local integration by parts, using that $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$, the identity $\langle\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}=-\langle{\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}},\frac{\mathrm{d}\boldsymbol{u}_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\rangle_{\Gamma_{\mathrm{D}}}$, and the Cauchy–Schwarz inequality, we obtain
@@
(6.6) $\begin{array}[]{ll}\displaystyle\mathcal{R}_{1}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))=-\sum_{K\in\mathcal{T}_{h}}\int_{K}\underline{\mathbf{curl}\,}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h}):({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})\\ \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Omega)}\int_{E}[\![(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}]\!]\cdot({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))\\ \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\int_{E}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h}-\boldsymbol{\nabla}\boldsymbol{u}_{\mathrm{D}})\boldsymbol{s}\cdot({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})\\ \qquad\displaystyle\leq\sum_{K\in\mathcal{T}_{h}}\|\underline{\mathbf{curl}\,}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\|_{0,K}\|{\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}\|_{0,K}\\ \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Omega)}\|[\![(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}]\!]\|_{0,E}\|{\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}\|_{0,E}+\gamma\|p-p_{h}\|_{0,\Omega}\,\|\mathbf{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}})\|_{0,\Omega}\\ \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\|(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}-\dfrac{\mathrm{d}\boldsymbol{u}_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\|_{0,E}\|{\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}\|_{0,E}.\end{array}$
@@
As in the previous result, the approximation properties of the Clément interpolation (see, e.g., [3, Section 3]) in conjunction with the Cauchy–Schwarz inequality, produces the desired result.
@@
Assume that there exists a convex domain $B$ such that $\Omega\subseteq B$ and $\Gamma_{\mathrm{N}}\subseteq\partial B$, and that $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$. Then, there exists a constant $C_{5}>0$, independent of $h$, such that
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0,\Omega}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,\Omega}\,\leq\,C_{5}\Bigg{\{}\,\sum_{T\in\mathcal{T}_{h}}\Xi_{s,K}^{2}\Bigg{\}}^{1/2}+C_{5}\gamma\|p-p_{h}\|_{0,\Omega}.$
Proof 6.10.
[…]
(6.7) $\boldsymbol{\tau}=\mbox{\boldmath{$\xi$}}+\mathbf{curl}\,{\boldsymbol{\chi}}{\quad\hbox{and}\quad}\|\mbox{\boldmath{$\xi$}}\|_{1,\Omega}+\|{\boldsymbol{\chi}}\|_{1,\Omega}\leq C_{\mathrm{Helm}}\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega},$
Using that $\mathcal{R}_{1}(\boldsymbol{\tau}_{h})=c(\boldsymbol{\tau}_{h},p_{h}-p)$ for all $\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}$, we have
$\mathcal{R}_{1}(\boldsymbol{\tau})=\mathcal{R}_{1}(\boldsymbol{\tau}-\boldsymbol{\tau}_{h})+c(\boldsymbol{\tau}_{h},p-p_{h})\quad\forall\,\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}.$
In particular, this holds for $\boldsymbol{\tau}_{h}$ defined as $\boldsymbol{\tau}_{h}=\mbox{\boldmath{$\Pi$}}_{h}\mbox{\boldmath{$\xi$}}+\mathbf{curl}\,(\mathbf{I}_{h}{\boldsymbol{\chi}})$, whence
$\mathcal{R}_{1}(\boldsymbol{\tau})=\mathcal{R}_{1}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}\mbox{\boldmath{$\xi$}})+\mathcal{R}_{1}(\mathrm{curl}\,({\boldsymbol{\chi}}-\mathbf{I}_{h}{\boldsymbol{\chi}}))+c(\boldsymbol{\tau}_{h},p-p_{h}).$
Hence, the proof follows from Lemmas 6.1, 6.5 and 6.7, and estimate (6.7).
@@
$\|{\boldsymbol{\varphi}}-{\boldsymbol{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}\,\leq\,C_{8}\Bigg{\{}\,\sum_{T\in\mathcal{T}_{h}}\Xi_{f,K}^{2}\Bigg{\}}^{1/2}+C_{8}\big{(}\|{\boldsymbol{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}+\gamma\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,\Omega}\big{)}.$
Proof 6.12.
@@
(6.8) $\begin{array}[]{ll}|\mathcal{R}_{2}(\mbox{\boldmath{$z$}}-\Pi_{h}(\mbox{\boldmath{$z$}}))|\,\leq C_{6}\,\|{\boldsymbol{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}\|\mbox{\boldmath{$z$}}\|_{1,\Omega}\\ \qquad\displaystyle+C_{6}\biggr{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h}-\nabla p_{h}\|_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|p_{\mathrm{D}}-p_{h}\|_{0,E}^{2}\biggr{)}^{1/2}\|\mbox{\boldmath{$z$}}\|_{1,\Omega},\\[17.22217pt] \displaystyle|\mathcal{R}_{2}(\mathrm{curl}\,(\phi-I_{h}\phi))|\leq C_{7}\|{\boldsymbol{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}\|\phi\|_{1,\Omega}\\[8.61108pt] \qquad\displaystyle+C_{7}\biggr{(}\,\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|\mathrm{rot}\,(\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h})\|_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Omega)}h_{E}\|[\![(\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h})\cdot\boldsymbol{s}]\!]\|_{0,E}^{2}\\[8.61108pt] \qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h}\cdot\boldsymbol{s}-\frac{\mathrm{d}p_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\|_{0,E}^{2}\biggr{)}^{1/2}\|\phi\|_{1,\Omega}.\end{array}$
Then, noting that $\mathcal{R}_{2}({\boldsymbol{\psi}}_{h})=0$ for all ${\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}$, and defining ${\boldsymbol{\psi}}_{h}$ as ${\boldsymbol{\psi}}_{h}=\Pi_{h}\mbox{\boldmath{$z$}}+\mathrm{curl}\,(I_{h}\phi)$, we have
$\mathcal{R}_{2}({\boldsymbol{\psi}})=\mathcal{R}_{2}({\boldsymbol{\psi}}-{\boldsymbol{\psi}}_{h})=\mathcal{R}_{2}(\mbox{\boldmath{$z$}}-\Pi_{h}\mbox{\boldmath{$z$}})+\mathcal{R}_{2}(\mathrm{curl}\,(\phi-I_{h}\phi)).$
Hence, the proof follows from Lemma 6.3, estimates (6.8), and the Helmholtz decomposition of $\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$.
@@
Assume that the hypotheses stated in Theorem 5.5 and Lemmas 6.9-6.11 are satisfied. Let $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (2.2) and (4.6), respectively. Assume further that
(6.9) $(C_{5}+C_{8})\gamma+C_{8}C^{*}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,)\leq\frac{1}{2}.$

### v2

@@ 6.1 Reliability of the a posteriori error estimator
$\begin{array}[]{ll}\displaystyle|\mathcal{R}_{1}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))|\leq C_{4}\gamma\|p-p_{h}||_{0,\Omega}\|{\bm{\chi}}\|_{1,\Omega}\\[8.61108pt] \displaystyle+C_{4}\biggr{(}\,\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}||\underline{\mathbf{curl}}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})||_{0,K}^{2}\\[8.61108pt] \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Omega)}h_{E}||[\![(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}]\!]||_{0,E}^{2}\\[8.61108pt] \displaystyle\qquad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}||(C^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}-\dfrac{\mathrm{d}{\bm{u}}_{\mathrm{D}}}{\mathrm{d}\bm{s}}||_{0,E}^{2}\biggr{)}^{1/2}\|{\bm{\chi}}\|_{1,\Omega}.\end{array}$
Proof.
Similarly to Lemma 6.3, adding ±$\pm\,c\,(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}),p_{h})$, we have
@@
$\begin{array}[]{ll}\mathcal{R}_{1}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))\\[4.30554pt] \quad\displaystyle=H(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))-c(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}),p)-a(\bm{\sigma}_{h},\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))\\[4.30554pt] \qquad\displaystyle-b(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}),({\bm{u}}_{h},\bm{\rho}_{h}))\\[4.30554pt] \displaystyle\quad=\langle(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))\\[8.61108pt] \displaystyle\quad-\int_{\Omega}\mathcal{C}^{-1}\bm{\sigma}_{h}:\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})-\int_{\Omega}\bm{\rho}_{h}:\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p_{h}\,\mathop{\rm tr}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})).\end{array}$
@@
Then, applying a local integration by parts, using that ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$, the identity $\langle\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}=-\langle{\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}},\frac{\mathrm{d}{\bm{u}}_{\mathrm{D}}}{\mathrm{d}\bm{s}}\rangle_{\Gamma_{\mathrm{D}}}$, and the Cauchy–Schwarz inequality, we obtain
@@
$\displaystyle\mathcal{R}_{1}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))$ $=-\sum_{K\in\mathcal{T}_{h}}\int_{K}\underline{\mathbf{curl}\,}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h}):({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})$
$\qquad+\sum_{E\in\mathcal{E}_{h}(\Omega)}\int_{E}[\![(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}]\!]\cdot({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})$
$\qquad-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))$
$\qquad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\int_{E}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h}-\bm{\nabla}{\bm{u}}_{\mathrm{D}})\bm{s}\cdot({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})$
$\displaystyle\leq\sum_{K\in\mathcal{T}_{h}}\|\underline{\mathbf{curl}\,}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\|_{0,K}\|{\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}\|_{0,K}$
$\qquad+\sum_{E\in\mathcal{E}_{h}(\Omega)}\|[\![(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}]\!]\|_{0,E}\|{\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}\|_{0,E}$
$\qquad\displaystyle+\gamma\|p-p_{h}\|_{0,\Omega}\,\|\mathbf{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}})\|_{0,\Omega}$
$\qquad+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\|(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}-\dfrac{\mathrm{d}{\bm{u}}_{\mathrm{D}}}{\mathrm{d}\bm{s}}\|_{0,E}\|{\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}\|_{0,E}.$
@@
As in the previous result, the approximation properties of the Clément interpolation (see, e.g., [3, Section 3]) in conjunction with the Cauchy–Schwarz inequality, produces the desired result. ∎
@@
Assume that there exists a convex domain $B$ such that $\Omega\subseteq B$ and $\Gamma_{\mathrm{N}}\subseteq\partial B$, and that ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$. Then, there exists a constant $C_{5}>0$, independent of $h$, such that
$\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|{\bm{u}}-{\bm{u}}_{h}\|_{0,\Omega}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,\Omega}\,\leq\,C_{5}\Bigg{\{}\,\sum_{T\in\mathcal{T}_{h}}\Xi_{s,K}^{2}\Bigg{\}}^{1/2}+C_{5}\gamma\|p-p_{h}\|_{0,\Omega}.$
Proof.
Let $\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$. From [3, Lemma 3.9], there exist $\bm{\xi}\in\mathbb{H}^{1}(\Omega)$ and ${\bm{\chi}}\in\mathbf{H}_{\Gamma_{\mathrm{N}}}^{1}(\Omega)$, such that
$\bm{\tau}=\bm{\xi}+\mathbf{curl}\,{\bm{\chi}}{\quad\hbox{and}\quad}\|\bm{\xi}\|_{1,\Omega}+\|{\bm{\chi}}\|_{1,\Omega}\leq C_{\mathrm{Helm}}\|\bm{\tau}\|_{\mathbf{div};\Omega},$ (6.4)
Using that $\mathcal{R}_{1}(\bm{\tau}_{h})=c(\bm{\tau}_{h},p_{h}-p)$ for all $\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}$, we have
$\mathcal{R}_{1}(\bm{\tau})=\mathcal{R}_{1}(\bm{\tau}-\bm{\tau}_{h})+c(\bm{\tau}_{h},p-p_{h})\quad\forall\,\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}.$
In particular, this holds for $\bm{\tau}_{h}$ defined as $\bm{\tau}_{h}=\mbox{\boldmath{$\Pi$}}_{h}\bm{\xi}+\mathbf{curl}\,(\mathbf{I}_{h}{\bm{\chi}})$, whence
$\mathcal{R}_{1}(\bm{\tau})=\mathcal{R}_{1}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}\bm{\xi})+\mathcal{R}_{1}(\mathrm{curl}\,({\bm{\chi}}-\mathbf{I}_{h}{\bm{\chi}}))+c(\bm{\tau}_{h},p-p_{h}).$
Hence, the proof follows from Lemmas 6.1, 6.3 and 6.4, and estimate (6.4). ∎
@@
$\|{\bm{\varphi}}-{\bm{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}\,\leq\,C_{8}\Bigg{\{}\,\sum_{T\in\mathcal{T}_{h}}\Xi_{f,K}^{2}\Bigg{\}}^{1/2}+C_{8}\big{(}\|{\bm{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}+\gamma\|\bm{\sigma}-\bm{\sigma}_{h}\|_{0,\Omega}\big{)}.$
Proof.
@@
$\begin{array}[]{ll}|\mathcal{R}_{2}(\mbox{\boldmath{$z$}}-\Pi_{h}(\mbox{\boldmath{$z$}}))|\,\leq C_{6}\,\|{\bm{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}\|\mbox{\boldmath{$z$}}\|_{1,\Omega}\\ \qquad\displaystyle+C_{6}\biggr{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h}-\nabla p_{h}\|_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|p_{\mathrm{D}}-p_{h}\|_{0,E}^{2}\biggr{)}^{1/2}\|\mbox{\boldmath{$z$}}\|_{1,\Omega},\\[17.22217pt] \displaystyle|\mathcal{R}_{2}(\mathrm{curl}\,(\phi-I_{h}\phi))|\leq C_{7}\|{\bm{\varphi}}_{h}\|_{0,\Omega}\|p-p_{h}\|_{0,\Omega}\|\phi\|_{1,\Omega}\\[8.61108pt] \qquad\displaystyle+C_{7}\biggr{(}\,\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}\|\mathrm{rot}\,({\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h})\|_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Omega)}h_{E}\|[\![({\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h})\cdot\bm{s}]\!]\|_{0,E}^{2}\\[8.61108pt] \qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h}\cdot\bm{s}-\frac{\mathrm{d}p_{\mathrm{D}}}{\mathrm{d}\bm{s}}\|_{0,E}^{2}\biggr{)}^{1/2}\|\phi\|_{1,\Omega}.\end{array}$ (6.5)
Then, noting that $\mathcal{R}_{2}({\bm{\psi}}_{h})=0$ for all ${\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h}$, and defining ${\bm{\psi}}_{h}$ as ${\bm{\psi}}_{h}=\Pi_{h}\mbox{\boldmath{$z$}}+\mathrm{curl}\,(I_{h}\phi)$, we have
$\mathcal{R}_{2}({\bm{\psi}})=\mathcal{R}_{2}({\bm{\psi}}-{\bm{\psi}}_{h})=\mathcal{R}_{2}(\mbox{\boldmath{$z$}}-\Pi_{h}\mbox{\boldmath{$z$}})+\mathcal{R}_{2}(\mathrm{curl}\,(\phi-I_{h}\phi)).$
Hence, the proof follows from Lemma 6.2, estimates (6.5), and the Helmholtz decomposition of $\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$. ∎
@@
Assume the hypotheses stated in Theorem 5.3 and Lemmas 6.5-6.6. Let $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (2.2) and (4.5), respectively. Assume further that
$(C_{5}+C_{8})\gamma+C_{8}C^{*}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,)\leq\frac{1}{2}.$ (6.6)